DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, January 29, 2004 Padelford C-36 1:30 pm

Uniqueness along a line for an inverse wave source problem

Jin Cheng (Fudan University, China)

In this talk, we will present an uniqueness result for an inverse wave source problem. We will show that partial information on the solution can determine partial information of the source function.

This is joint work with G. H. Ding (Fudan University) and M. Yamamoto (University of Tokyo).

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.